Evaluate the following integral after converting to polar coordinates: ∫(0 to 1)∫(y to sqrt(2-y^2)) (x+y) dx dy
1
Expert's answer
2020-11-08T18:35:24-0500
"Solution"

STEP 1

Convert the upper bound of the integral for "x" into polar coordinates. The result is a circle of radius "\sqrt 2" centered at the origin.


"x=\sqrt{2-y^2} \implies x^2=2-y^2 \implies x^2+y^2=2\\
r^2=2 \implies r=\sqrt 2"

STEP 2

Convert the other bounds into polar coordinates. These tell us that the angle ranges from to "\frac{\pi}{4}"

I.e the region is a wedge of the circle of radius "\sqrt 2", with an angle sweeping from zero to "\frac{\pi}{4}"

"x=y \implies \theta=\frac{\pi}{4}\\
y=0 \implies \theta=0"

STEP 3

Convert the function into polar coordinates


"=x+y = r \cdot cos \theta + r \cdot sin \theta\\
=\intop_0^{\frac{\pi}{4}} \intop_0^{\sqrt{2}}( r \cdot cos \theta + r \cdot sin \theta)r\ \delta r\ \delta \theta\\
=\intop_0^{\frac{\pi}{4}} \intop_0^{\sqrt{2}}r^2( cos \theta + sin \theta)r\ \delta r\ \delta \theta\\"

STEP 4

Integrate the function with respect to "r"


"= \intop_0 ^{\frac{\pi}{4}}[\frac{1}{3}r^3(cos \theta+sin \theta)]_0^{\sqrt{2}} \delta \theta\\
= \intop_0 ^{\frac{\pi}{4}}\frac{2 \sqrt{2}}{3}(cos \theta+sin \theta) \delta \theta"

STEP 5

Integrate the function with respect to theta

"= [\frac{2 \sqrt{2}}{3}(sin \theta-cos \theta) ]_0^{\frac{\pi}{4}}=\frac{2 \sqrt{2}}{3}---->Answer"